# on dextral symmetric algebras

### v1

@@ Abstract.
dextral symmetric, symmetric rings, Leibniz algebra, Left nilpotent right Leibniz algebra, solvable Leibniz algebra, Leavitt path algebra,
@@ 1. Introduction
Symmetric rings introduced by Lambek in [1] in the class of associative rings with unity undoubtedly plays an important role in the domain of ring and module theory. Along with the extensive studies that have been conducted on the symmetric rings, algebraists have also introduced many generalizations of these rings, [13, 14, 15, 16]. This motivated us to extend this notion to various algebras (not necessarily associative) by introducing the notion of dextral symmetric algebras.
@@
Recall that an ideal $I$ of an associative ring $R$ is called symmetric if for any $x,y,z\in R$, $xyz\in I$ implies $xzy\in I$. Lambek called an associative ring with unity symmetric if the zero ideal is symmetric. We generalize the definition of symmetricity of an ideal of a ring to an ideal $I$ of an algebra $A$ by calling it a symmetric ideal if it has the property that if $a(bc)\in I$, then $b(ac)\in I$, for any $a,b,c\in A$. We call an algebra $A$ (not nesessarily associative) a dextral symmetric algebra if the zero ideal of $A$ is symmetric. Further, an ideal of an algebra is called a dextral symmetric ideal if it is dextral symmetric as a subalgebra. It is worth mentioning here that in associative algebras with unity both the symmetric and dextral symmetric conditions are equivalent.
@@
The results on symmetric rings intrigued us to explore if those results are true in the broader picture as well. For example, in a symmetric ring $R$ we have, $r_{1}r_{2}\dots r_{n}=0$ implies $r_{\sigma(1)}r_{\sigma(2)}\dots r_{\sigma(n)}=0$, where $\sigma$ is any permutation of the set $\{1,2,\dots,n\}$, $r_{i}\in R$ and, $n\in\mathbb{N}$, [13, Lemma 1.2]. Although this result may not be readily true in the case of a non-associative dextral symmetric algebra, in Proposition 4.1 and Theorem 4.2, we observe a close result for dextral symmetric Leibniz algebras. This further prompted us to introduce another subclass of the class of right Leibniz algebras called the class of left nilpotent right Leibniz algebra. We show that there is a close connection between solvability, dextral symmetricity and left nilpotency within the class of right Leibniz algebra.
@@
Another observation is that, in the class of anti-commutative algebras the dextral symmetric algebras are in fact equivalent to the CB-algebras introduced by Saha and Towers in [4]. Recall here that an algebra $A$ is called anti-commutative if for each $x\in A$, $x^{2}=0$. This also implies that in such an algebra $A$, all $x,y\in A$ satisfy $xy=-yx$. We remind here that an algebra $A$ is called a CB-algebra (as in [4]) if whenever $xy=0$ for any $x,y\in A$, we have $(xz)y=0$ for all $z\in A$. In anti-commuatative algebras, the CB-algebras have an important property that every centralizer of these algebras is an ideal (See [4]).
@@
In Section 2, we document some examples of dextral symmetric algebras in the context of (anti) commutative and (anti) associative algebras. As mentioned in the previous paragraph, we show here that in anti-commutative algebras, the notions of dextral symmetric algebras, anti-associative algebras as well as the CB-algebras are equivalent. While in Section 3, we consider the class of Leavitt path algebras and determine the graphs over which a Leavitt path algebra over a commutative ring with unity will be dextral symmetric which leads to the complete classification of a dextral symmetric Leavitt path algebras. Note here that a Leavitt path algebra of a graph over a commutative ring with unity is associative. Section 4 focuses on results on right Leibniz algebras (which are non-commutative, non-associative algebras) that are dextral symmetric. We also introduce a weaker form of nilpotency for a right Leibniz algebra namely left nilpotency and establish an equivalency between a solvable right Leibniz algebra and a left nilpotent right Leibniz algebra through the dextral symmetricity. Finally, in Section 5 we classify the solvable dextral symmetric right Leibniz algebras upto dimension $4$.
@@ 2. Dextral symmetric algebras
In this section we state some basic properties of dextral symmetric algebras and provide some examples of the dextral symmetric algebras based on commuatativity and associativity of the algebra. We also show that in the class of anti-commutative algebras, the classes of dextral symmetric, anti-associative and CB algebras coincide.
@@
Every subalgebra of a dextral symmetric algebra is dextral symmetric. Moreover, dextral symmetric algebras are preserved under finite direct sums. Thus, every algebra has a largest dextral symmetric ideal.
@@
Proposition 2.5.
Let $A_{1},A_{2}$ be algebras, and $\phi:A_{1}\to A_{2}$ be an injective homomorphism. If $A_{1}$ is dextral symmetric then so is $\phi(A_{1})$.
@@
We cannot relax the injective hypothesis of Proposition 2.5, as can be seen by taking the natural projection from the free algebra $A$ of Example 2.3 onto the algebra $A/I$.
@@ 3. Classification of dextral symmetric Leavitt path algebras
In this section we consider a Leavitt path algebra over a unital commutative ring $R$. We will first characterize the graph associated to a Leavitt path alegbra that is dextral symmetric. For which we will start with mentioning some basic definitions concerning this class of algebras which will be helpful to us in establishing our aim. Finally we will give a complete classification of a dextral symmetric Leavitt path algebra whose coefficients are taken over a unital commutative ring.
By referring to a graph in this paper, we mean a directed graph $E:=(E^{0},E^{1},r,s),$ where $E^{0}$ and $E^{1}$ are countable sets of vertices and edges respectively. The map $r,s:E^{1}\rightarrow E^{0}$ carries each edge to its range and source respectively.
A path in a graph $E$ of length $n$ is a sequence of $n$ edges and is given by $\beta:=e_{1}e_{2}\dots e_{n}$, where $r(e_{i})=s(e_{i+1})$ for $i\in\{1,\dots n-1\}.$ The vertices in $E^{0}$ are considered to be paths of length 00. We denote the set of all finite paths by $E^{*}:=\cup_{n=0}^{\infty}E^{n}$, where $E^{n}$ denotes the set of all paths of length $n$.
For an edge $e\in E^{1}$ with $s(e)=u,~{}r(e)=v$, where $u,v\in E^{0}$, we have a corresponding ghost edge $e^{*}$ such that $s(e^{*})=v,~{}r(e^{*})=u.$ We denote the collection of all such ghost edges by $(E^{1})^{*}$. A loop is an edge $e\in E^{1}$ such that $s(e)=r(e)$.
@@
The Leavitt path algebra of an arbitrary directed graph $E$ over a unital commutative ring $R$, denoted by $L_{R}(E)$ is an $R$-algebra generated by a set of pair-wise orthogonal idempotent elements $\{v~{}|~{}v\in E^{0}\}$ and a set $\{e,e^{*}~{}|~{}e\in E^{*}\}$ of edges and ghost edges of $E$ satisfying the following conditions:
@@
Let us prove the lemma by contradiction. Let $L_{R}(E)$ be a dextral symmetric Leavitt path algebra. If possible, assume that the graph $E$ contains an edge $e\in E^{1}$ such that $s(e)\neq r(e)$. Then we have $es(e)r(e)=0$, however $s(e)er(e)=e\neq 0$ (Remark 3.1). This is a contradiction to our assumption that $L_{R}(E)$ is dextral symmetric. Hence $E$ contains no edge with distinct source and range. ∎
@@
The proof is easy to see as the Leavitt path algebra of such a graph is isomorphic to the Laurent polynomial algebra $R[x,x^{-1}]$ for an indeterminate $x$ [3, Proposition 1.3.4]. Since $R$ is commutative, $R[x,x^{-1}]$ is commutative. Therefore, $a(bc)=0\implies b(ac)=0$ for all $a,b,c\in R[x,x^{-1}]$ yielding that it is dextral symmetric. ∎
@@
A dextral symmetric Leavitt path alegebra $L_{R}(E)$ is isomorphic to:
@@
for some index sets $I$, $J$.
@@ 4. Leibniz algebra
$[a,[b,c]]=[[a,b],c]+[b,[a,c]]~{}\text{for each $a,b,c\in\mathscr{L}$}.$
We start off our study of dextral symmetric right Leibniz algebra with the next proposition. We show that for every triplet of elements $x,y,z$ in a dextral symmetric right Leibniz algebra, the product of the form $[x,[y,z]]$ is equal to the product (of the same form) obtained by any permutation of $x,y$ and $z$ up to sign. Specifically, if the permutation is taken clockwise the sign is positive otherwise the sign is negative.
@@
If a right Leibniz algebra $\mathscr{L}$ is dextral symmetric, then for any $x,y,z\in\mathscr{L}$, the following relations hold:
@@
Since $\mathscr{L}$ is a right Leibniz algebra, we always have $[y,[x+z,x+z]]=0$ for all $x,y,z\in\mathscr{L}$. Moreover, $\mathscr{L}$ is dextral symmetric and hence we have,
$[x+z,[y,x+z]]=0\implies[x,[y,z]]=-[z,[y,x]].$
It also follows from the Leibniz identity that $[x,[y,z]]=-[x,[z,y]].$ Now replacing $x$ by $y$, $y$ by $z$, and $z$ by $x$, the same argument can be applied to get the required result. ∎
@@
Over the years many studies have been conducted on solvable and right nilpotent Leibniz algebras [5, 7, 10, 11, 12]. Here, we introduce the notion of a left nilpotent Leibnizalgebra which is weaker than the usual notion of nilpotency. However, these two notions of nilpotency coincide for a Lie algebra. In order to give the formal definition of this new class of Leibniz algebras, let us mention the following four sequences of two-sided ideals of a given Leibniz algebra $\mathscr{L}$.
@@
The algebra $\mathscr{L}$ is said to be nilpotent if $\mathscr{L}^{n}=0$, for some $n\in\mathbb{N}$, right nilpotent if $\mathscr{L}^{<k>}=0$, for some $k\in\mathbb{N}$, and solvable if there exists $n^{\prime}\in\mathbb{N}$ such that $\mathscr{L}^{[n^{\prime}]}=0$, [5].
@@
It is easy to see from the Leibniz identity that in a Leibniz algebra, right nilpotency will imply left nilpotency. But Example 4.3 below shows that the reverse implication may not hold in general. Thus implying that in a Leibniz algebra, right nilpotency and left nilpotency are two different concepts. This distinction comes mainly from the fact that the brackets in a generic Leibniz algebra do not commute. .
@@
Consider the 3-dimensional Leibniz algebra, $\mathscr{L}^{\prime}$ generated by $\{x,y,z\}$ with the non-zero products given by:
@@
Thus, in particular for any integers $m_{1},m_{2}\geq 1$, and $n\geq 2$,
@@
Lemma 4.6 shows that for a dextral symmetric Leibniz algebra $\mathscr{L},~{}\mathscr{L}^{[3]}=\mathscr{L}^{(4)}$. We will generalize this observation in the subsequent proposition.
@@
    Since in a right Leibniz algebra which is also a left Leibniz algebra, the concepts of left nilpotency and right nilpotency are the same, it follows by Theorem 4.11 that in a dextral symmetric right Leibniz algebra which is also a left Leibniz algebra all three notions of right nilpotency, left nilpotency and solvability coincide.
@@ 5. Classification of complex dextral symmetric Leibniz algebras upto dimension $4$
In this section, we consider the Leibniz algebras over the field of complex numbers. It follows from Corollary 4.14 that to classify $4$-dimensional dextral symmetric Leibniz algebra, it is enough to classify them based on the classification of $4$-dimensional solvable Leibniz algebras in [7] and [8].
@@
If $\mathscr{L}$ is a $2$-dimensional right nilpotent right Liebniz algebra, then $\mathscr{L}$ is dextral symmetric.
@@
There are exactly $17$ isomorphism classes of $4$-dimensional right nilpotent dextral symmetric Leibniz algebras.
@@
According to [7], the isomorphism classes of right nilpotent Leibniz algebras are given by the following algebras:
$\mathscr{Q}_{1}$ $:~{}[x,x]=y,~{}[y,x]=z,~{}[z,x]=w~{};$
$\mathscr{Q}_{2}$ $:~{}[x,x]=z,~{}[x,y]=w,~{}[y,x]=z,~{}[z,x]=w~{};$
$\mathscr{Q}_{3}$ $:~{}[x,x]=z,~{}[y,x]=z,~{}[z,x]=w~{};$
$\mathscr{Q}_{4}(\alpha)$ $:~{}[x,x]=z,~{}[x,y]=\alpha w,~{}[y,x]=z,~{}[y,y]=w,~{}[z,x]=w,~{}\alpha\in\{0,1\}~{};$
$\mathscr{Q}_{5}$ $:~{}[x,x]=z,~{}[x,y]=w,~{}[z,x]=w~{};$
$\mathscr{Q}_{6}$ $:~{}[x,x]=z,~{}[y,y]=w,~{}[z,x]=w~{};$
$\mathscr{Q}_{7}$ $:~{}[x,x]=w,~{}[y,x]=z,~{}[z,x]=w,~{}[x,y]=-z,~{}[x,z]=-w~{};$
$\mathscr{Q}_{8}$ $:~{}[x,x]=w,~{}[y,x]=z,~{}[z,x]=w,~{}[x,y]=-z+w,~{}[x,z]=-w~{};$
$\mathscr{Q}_{9}$ $:~{}[x,x]=w,~{}[y,x]=z,~{}[y,y]=w,~{}[z,x]=w,~{}[x,y]=-z+2w,~{}[x,z]=-w~{};$
$\mathscr{Q}_{10}$ $:~{}[x,x]=w,~{}[y,x]=z,~{}[y,y]=w,~{}[z,x]=w,~{}[x,y]=-z,~{}[x,z]=-w~{};$
$\mathscr{Q}_{11}$ $:~{}[x,x]=w,~{}[x,y]=z,~{}[y,x]=-z,~{}[y,y]=-2z+w~{};$
$\mathscr{Q}_{12}$ $:~{}[x,y]=z,~{}[y,x]=w,~{}[y,y]=-z~{};$
$\mathscr{Q}_{13}(\alpha)$ $:~{}[x,x]=z,~{}[x,y]=w,~{}[y,x]=-\alpha z,~{}[y,y]=-w,~{}\alpha\in\mathbb{C}~{};$
$\mathscr{Q}_{14}(\alpha)$ $:~{}[x,x]=w,~{}[x,y]=\alpha w,~{}[y,x]=-\alpha w,~{}[y,y]=w,~{}[z,z]=w,~{}\alpha\in\mathbb{C}~{};$
$\mathscr{Q}_{15}$ $:~{}[x,y]=w,~{}[x,z]=w,~{}[y,x]=-w,~{}[y,y]=w,~{}[z,x]=w~{}:$
$\mathscr{Q}_{16}$ $:~{}[x,x]=w,~{}[x,y]=w,~{}[y,x]=-w,~{}[z,z]=w~{};$
$\mathscr{Q}_{17}$ $:~{}[x,y]=z,~{}[y,x]=w~{};$
$\mathscr{Q}_{18}$ $:~{}[x,y]=z,~{}[y,x]-z,~{}[y,y]=w~{};$
$\mathscr{Q}_{19}$ $:~{}[y,x]=w,~{}[y,y]=z~{};$
$\mathscr{Q}_{20}(\alpha)$ $:~{}[x,y]=w,~{}[y,x]=\frac{1+\alpha}{1-\alpha}w,~{}[y,y]=z,~{}\alpha\in\mathbb{C}\setminus\{1\}~{};$
$\mathscr{Q}_{21}$ $:~{}[x,y]=w,~{}[y,x]=-w,~{}[z,z]=w.$
Notice that in each of the algebras from $\mathscr{Q}_{7}$ to $\mathscr{Q}_{10}$ we have $[y,[x,x]]=0$ but $[x,[y,x]]\neq 0$. Whereas, for all the other $17$ isomorphism classes, $\mathscr{Q}_{i}^{(3)}=0$. ∎
@@
Assume for the sake of contradiction that $\mathscr{S}_{2}$ is dextral symmetric. Then $[b,[a,b]]=0$ for all $a,b\in\mathscr{S}_{2}$. However, $[x,[y,x]]=-y\neq 0$. Thus $\mathscr{S}_{2}$ can not be dextral symmetric. ∎
@@
$\underbrace{[w,[w,[w,\cdots w,[w}_{n-times},y]]]]=(-1)^{n}y\neq 0,\text{ for any }n\in\mathbb{N}.$
@@
Any solvable Leibniz algebra with dimension $4$ and nilradical isomorphic to $\mu_{3}$ is isomorphic to one of the following algebras which are pairwise isomorphic:
@@
Similar to the proof of Theorem 5.4, we can show that in each of the above mentioned algebras, there exists elements $a,b$ such that $[b,[a,b]]\neq 0$. Which yields that $\mathscr{L}_{1}(\lambda)$, $\mathscr{L}_{2}$ and $\mathscr{L}_{3}$ are not dextral symmetric. ∎

### v2

@@ Abstract.
dextral symmetric, symmetric rings, Leibniz algebra, Left nilpotent right Leibniz algebra, solvable Leibniz algebra, Leavitt path algebra
@@ 1. Introduction
Symmetric rings, introduced by Lambek in [1] as a subclass of the class of associative rings with unity, undoubtedly play an important role in the domain of ring and module theory. Along with the extensive studies that have been conducted on symmetric rings, algebraists have also introduced many generalizations of these rings (see, for example, [14, 15, 16, 17]). This motivated us to extend this notion to various algebras (not necessarily associative) by introducing the notion of dextral symmetric algebras.
@@
Recall that an ideal $I$ of an associative ring $R$ is called symmetric if for any $x,y,z\in R$, $xyz\in I$ implies $xzy\in I$. Lambek called an associative ring with unity symmetric if the zero ideal is symmetric. We generalize the definition of symmetricity of an ideal of a ring to an ideal $I$ of an algebra $A$ by calling it a symmetric ideal if it has the property that if $a(bc)\in I$, then $b(ac)\in I$, for any $a,b,c\in A$. We call an algebra $A$ (not necessarily associative) a dextral symmetric algebra if the zero ideal of $A$ is symmetric. Further, an ideal of an algebra is called a dextral symmetric ideal if it is dextral symmetric as a subalgebra. It is worth mentioning here that in associative algebras with unity, both the symmetric and dextral symmetric conditions are equivalent.
@@
The results on symmetric rings intrigued us to explore whether those results hold in a broader context as well. For example, in a symmetric ring $R$, we have $r_{1}r_{2}\dots r_{n}=0$ implies $r_{\sigma(1)}r_{\sigma(2)}\dots r_{\sigma(n)}=0$, where $\sigma$ is any permutation of the set $\{1,2,\dots,n\}$, $r_{i}\in R$ and $n\in\mathbb{N}$ [14, Lemma 1.2]. Although this result may not be readily true in the case of a non-associative dextral symmetric algebra, in Proposition 4.1 and Theorem 4.2, we observe a similar result for dextral symmetric Leibniz algebras. This further prompted us to introduce another subclass of the class of right Leibniz algebras called the class of left nilpotent right Leibniz algebra. We show that there is a close connection between solvability, dextral symmetricity, and left nilpotency within the class of right Leibniz algebras.
@@
Another observation is that in the class of anti-commutative algebras, dextral symmetric algebras are, in fact, equivalent to the CB-algebras introduced by Saha and Towers in [4]. Recall here that an algebra $A$ is called anti-commutative if for each $x\in A$, $x^{2}=0$. This also implies that in such an algebra $A$, all $x,y\in A$ satisfy $xy=-yx$. We remind here that an algebra $A$ is called a CB-algebra (as in [4]) if whenever $xy=0$ for any $x,y\in A$, we have $(xz)y=0$ for all $z\in A$. In anti-commutative algebras, the CB-algebras display an interesting property: namely, that every centralizer of these algebras is an ideal (see [4]).
@@
In Section 2, we document some examples of dextral symmetric algebras in the context of (anti)commutative and (anti)associative algebras. As mentioned in the previous paragraph, we show that in anti-commutative algebras, the notions of dextral symmetric algebras, anti-associative algebras, and CB-algebras are equivalent. In Section 3, we consider the class of Leavitt path algebras and determine the graphs over which a Leavitt path algebra over a commutative ring with unity will be dextral symmetric. This leads to the complete classification of dextral symmetric Leavitt path algebras. Note that a Leavitt path algebra of a graph over a commutative ring with unity is associative. Section 4 focuses on results concerning right Leibniz algebras (which are non-commutative, non-associative algebras) that are dextral symmetric. We also introduce a weaker form of nilpotency for a right Leibniz algebra, namely left nilpotency, and establish an equivalency between a solvable right Leibniz algebra and a left nilpotent right Leibniz algebra through dextral symmetricity. Finally, in Section 5, we classify the solvable dextral symmetric right Leibniz algebras up to dimension 4.
@@ 2. Dextral symmetric algebras
In this section, we state some basic properties of dextral symmetric algebras and provide examples based on the commutativity and associativity of the algebra. We also show that in the class of anti-commutative algebras, the classes of dextral symmetric, anti-associative, and CB algebras coincide.
@@
Every subalgebra of a dextral symmetric algebra is dextral symmetric. Moreover, dextral symmetric algebras are preserved under finite direct sums. Thus, every algebra has a largest dextral symmetric ideal. A homomorphic image of a dextral symmetric algebra under an injective homomorphism is also dextral symmetric.
@@
The injective hypothesis of remark 2.1 cannot be relaxed, as can be seen by taking the natural projection from the free algebra $A$ of Example 2.3 onto the algebra $A/I$.
@@ 3. Classification of dextral symmetric Leavitt path algebras
In this section we consider a Leavitt path algebra over a unital commutative ring $R$. We first characterize the graph associated to a Leavitt path algebra when it is dextral symmetric. We begin by mentioning some basic definitions concerning this class of algebras that are helpful to us in establishing our aim. Finally, we provide a complete classification of dextral symmetric Leavitt path algebras with coefficients taken from a unital commutative ring.
By referring to a graph in this paper, we mean a directed graph $E:=(E^{0},E^{1},r,s),$ where $E^{0}$ and $E^{1}$ are countable sets of vertices and edges respectively. The map $r,s:E^{1}\rightarrow E^{0}$ carries each edge to its range and source, respectively.
A loop is an edge $e\in E^{1}$ such that $s(e)=r(e)$. A path in a graph $E$ of length $n$ is a sequence of $n$ edges given by $\beta:=e_{1}e_{2}\dots e_{n}$, where $r(e_{i})=s(e_{i+1})$ for $i\in\{1,\dots n-1\}.$ The vertices in $E^{0}$ are considered to be paths of length 00. We denote the set of all finite paths by $E^{*}:=\cup_{n=0}^{\infty}E^{n}$, where $E^{n}$ denotes the set of all paths of length $n$.
For an edge $e\in E^{1}$ with $s(e)=u,~{}r(e)=v$, where $u,v\in E^{0}$, we have a corresponding ghost edge $e^{*}$ such that $s(e^{*})=v,~{}r(e^{*})=u.$ We denote the collection of all such ghost edges by $(E^{1})^{*}$.
@@
The Leavitt path algebra of an arbitrary directed graph $E$ over a unital commutative ring $R$, denoted by $L_{R}(E)$ is an $R$-algebra generated by a set of pairwise orthogonal idempotent elements $\{v~{}|~{}v\in E^{0}\}$ and a set $\{e,e^{*}~{}|~{}e\in E^{*}\}$ of edges and ghost edges of $E$ satisfying the following conditions :
@@
Let us prove the lemma by contradiction. Consider a dextral symmetric Leavitt path algebra $L_{R}(E)$. Suppose, that the graph $E$ contains an edge $e\in E^{1}$ such that $s(e)\neq r(e)$. Then we would have $es(e)r(e)=0$. However, $s(e)er(e)=e\neq 0$ (as mentioned in Remark 3.1). This contradicts our assumption that $L_{R}(E)$ is dextral symmetric. Thus, $E$ contains no edge with distinct source and range. ∎
@@
The proof is straightforward. The Leavitt path algebra of such a graph is isomorphic to the Laurent polynomial algebra $R[x,x^{-1}]$ for an indeterminate $x$, as shown in [3, Proposition 1.3.4]. Since $R$ is commutative, $R[x,x^{-1}]$ is commutative. Therefore, by part (i) of Proposition 2.6, $R[x,x^{-1}]$ is dextral symmetric. ∎
@@
A dextral symmetric Leavitt path algebra $L_{R}(E)$ is isomorphic to :
@@
for some index sets $I$ and $J$.
@@ 4. Leibniz algebra
[…]
We start off our study of dextral symmetric right Leibniz algebras with the next proposition. We show that for every triplet of elements $x,y,z$ in a dextral symmetric right Leibniz algebra, the product of the form $[x,[y,z]]$ is equal to the product (of the same form) obtained by any permutation of $x,y$ and $z$ up to sign. Specifically, if the permutation is taken clockwise the sign is positive otherwise the sign is negative.
@@
If a right Leibniz algebra $\mathscr{L}$ is dextral symmetric, then for all $x,y,z\in\mathscr{L}$, the following relations hold:
@@
Since $\mathscr{L}$ is a right Leibniz algebra, we always have $[y,[x+z,x+z]]=0$ for all $x,y,z\in\mathscr{L}$. Moreover, $\mathscr{L}$ is dextral symmetric, which implies that $[x+z,[y,x+z]]=0$, leading to $[x,[y,z]]=-[z,[y,x]]$. It follows from the Leibniz identity that $[x,[y,z]]=-[x,[z,y]]$. By replacing $x$ with $y$, $y$ with $z$, and $z$ with $x$, the same argument can be applied to obtain the required result. ∎
@@
Over the years, many studies have been carried out on solvable and right nilpotent Leibniz algebras, [5, 7, 10, 11, 12, 13]. In this paper, we introduce another class of Leibniz algebras that lies between the classes of right nilpotent and solvable Leibniz algebras. We introduce the notion of left nilpotency, which is a weaker notion than the usual notion of nilpotency. However, these two notions of nilpotency coincide for a Lie algebra. In order to give the formal definition of this new class of Leibniz algebras, let us mention the following four sequences of two-sided ideals of a given Leibniz algebra $\mathscr{L}$.
@@
The algebra $\mathscr{L}$ is said to be nilpotent if $\mathscr{L}^{n}=0$, for some $n\in\mathbb{N}$, right nilpotent if $\mathscr{L}^{<k>}=0$, for some $k\in\mathbb{N}$, and solvable if there exists $n^{\prime}\in\mathbb{N}$ such that $\mathscr{L}^{[n^{\prime}]}=0$ [5].
@@
It is easy to see from the Leibniz identity that in a Leibniz algebra, right nilpotency will imply left nilpotency. But Example 4.3 below shows that the reverse implication may not hold in general. Thus implying that in a Leibniz algebra, right nilpotency and left nilpotency are two different concepts. This distinction comes mainly from the fact that the brackets in a generic Leibniz algebra do not commute.
@@
Consider the 3-dimensional Leibniz algebra $\mathscr{L}^{\prime}$ generated by $\{x,y,z\}$ with the non-zero products given by:
@@
Thus, in particular for any three integers $m_{1},m_{2}\geq 1$, and $n\geq 2$,
@@
Lemma 4.6 shows that for a dextral symmetric Leibniz algebra $\mathscr{L},~{}\mathscr{L}^{[3]}=\mathscr{L}^{(4)}$. We generalize this observation in the subsequent propositions.
@@
    Since in a right Leibniz algebra which is also a left Leibniz algebra, the concepts of left nilpotency and right nilpotency are the same, it follows from Theorem 4.11 that in a dextral symmetric right Leibniz algebra which is also a left Leibniz algebra all three notions of right nilpotency, left nilpotency and solvability coincide.
@@ 5. Classification of complex dextral symmetric Leibniz algebras up to dimension 4
In this section, we consider the Leibniz algebras over the field of complex numbers. It follows from Corollary 4.14 that to classify 44-dimensional dextral symmetric Leibniz algebra, it is enough to classify them based on the classification of 44-dimensional solvable Leibniz algebras in [7], [13] and [8].
@@
If $\mathscr{L}$ is a $2$-dimensional right nilpotent Liebniz algebra, then $\mathscr{L}$ is dextral symmetric.
@@
There are exactly $18$ isomorphism classes of $4$-dimensional right nilpotent dextral symmetric Leibniz algebras.
@@
According to [7] and [13], the isomorphism classes of right nilpotent Leibniz algebras are given by the following algebras.
$\mathscr{N}_{1}$ $:~{}[x,x]=y,~{}[y,x]=z,~{}[z,x]=w~{};$
$\mathscr{N}_{2}$ $:~{}[x,x]=z,~{}[x,y]=w,~{}[y,x]=z,~{}[z,x]=w~{};$
$\mathscr{N}_{3}$ $:~{}[x,x]=z,~{}[y,x]=z,~{}[z,x]=w~{};$
$\mathscr{N}_{4}(\alpha)$ $:~{}[x,x]=z,~{}[x,y]=\alpha w,~{}[y,x]=z,~{}[y,y]=w,~{}[z,x]=w,~{}\alpha\in\{0,1\}~{};$
$\mathscr{N}_{5}$ $:~{}[x,x]=z,~{}[x,y]=w,~{}[z,x]=w~{};$
$\mathscr{N}_{6}$ $:~{}[x,x]=z,~{}[y,y]=w,~{}[z,x]=w~{};$
$\mathscr{N}_{7}$ $:~{}[x,x]=w,~{}[y,x]=z,~{}[z,x]=w,~{}[x,y]=-z,~{}[x,z]=-w~{};$
$\mathscr{N}_{8}$ $:~{}[x,x]=w,~{}[y,x]=z,~{}[z,x]=w,~{}[x,y]=-z+w,~{}[x,z]=-w~{};$
$\mathscr{N}_{9}$ $:~{}[x,x]=w,~{}[y,x]=z,~{}[y,y]=w,~{}[z,x]=w,~{}[x,y]=-z+2w,~{}[x,z]=-w~{};$
$\mathscr{N}_{10}$ $:~{}[x,x]=w,~{}[y,x]=z,~{}[y,y]=w,~{}[z,x]=w,~{}[x,y]=-z,~{}[x,z]=-w~{};$
$\mathscr{N}_{11}$ $:~{}[x,x]=w,~{}[x,y]=z,~{}[y,x]=-z,~{}[y,y]=-2z+w~{};$
$\mathscr{N}_{12}$ $:~{}[x,y]=z,~{}[y,x]=w,~{}[y,y]=-z~{};$
$\mathscr{N}_{13}(\alpha)$ $:~{}[x,x]=z,~{}[x,y]=w,~{}[y,x]=-\alpha z,~{}[y,y]=-w,~{}\alpha\in\mathbb{C}~{};$
$\mathscr{N}_{14}(\alpha)$ $:~{}[x,x]=w,~{}[x,y]=\alpha w,~{}[y,x]=-\alpha w,~{}[y,y]=w,~{}[z,z]=w,~{}\alpha\in\mathbb{C}~{};$
$\mathscr{N}_{15}$ $:~{}[x,y]=w,~{}[x,z]=w,~{}[y,x]=-w,~{}[y,y]=w,~{}[z,x]=w~{}:$
$\mathscr{N}_{16}$ $:~{}[x,x]=w,~{}[x,y]=w,~{}[y,x]=-w,~{}[z,z]=w~{};$
$\mathscr{N}_{17}$ $:~{}[x,y]=z,~{}[y,x]=w~{};$
$\mathscr{N}_{18}$ $:~{}[x,y]=z,~{}[y,x]-z,~{}[y,y]=w~{};$
$\mathscr{N}_{19}$ $:~{}[y,x]=w,~{}[y,y]=z~{};$
$\mathscr{N}_{20}(\alpha)$ $:~{}[x,y]=w,~{}[y,x]=\frac{1+\alpha}{1-\alpha}w,~{}[y,y]=z,~{}\alpha\in\mathbb{C}\setminus\{1\}~{};$
$\mathscr{N}_{21}$ $:~{}[x,y]=w,~{}[y,x]=-w,~{}[z,z]=w~{};$
$\mathscr{N}_{22}$ $:~{}[z,x]=w,~{}[y,z]=w.$
Notice that in each of the algebras from $\mathscr{N}_{7}$ to $\mathscr{N}_{10}$ we have $[y,[x,x]]=0$ but $[x,[y,x]]\neq 0$. Whereas, for all the other $18$ isomorphism classes, $\mathscr{N}_{i}^{(3)}=0$. ∎
@@
Assume for contradiction that $\mathscr{S}_{2}$ is dextral symmetric. Then $[b,[a,b]]=0$ for all $a,b\in\mathscr{S}_{2}$. However, $[z,[x,z]]=-x\neq 0$. Thus $\mathscr{S}_{2}$ can not be dextral symmetric. ∎
@@
$\underbrace{[w,[w,[w,\cdots w,[w}_{n-times},y]\cdots]]]=(-1)^{n}y\neq 0,\text{for any }n\in\mathbb{N}.$
@@
Any $4$-dimensional solvable Leibniz algebra with nilradical isomorphic to $\mu_{2}$ is isomorphic to one of the following algebras which are pairwise non-isomorphic:
@@
Similar to the proof of Theorem 5.4, we can show that in each of the above mentioned algebras, there exists elements $a,b$ such that $[b,[a,b]]\neq 0$. Which yields that the algebras $\mathscr{L}_{1}(\lambda)$, $\mathscr{L}_{2}$ and $\mathscr{L}_{3}$ are not dextral symmetric. ∎